13. f(x) = -13x2 - 7x4 - 11x
find f(5)
A. 538
B. -6288203
C. 639
D. -4595
E. -4,755
F. -3456

14. f(x) = 15x - 12
g(x) = -15x2 + 14x - 10
find g(g(7))
A. -6288203
B. 6543
C. 639
D. -4595
E. 538
F. -162,632

Found 2 solutions by math_tutor2020, josgarithmetic:
Answer by math_tutor2020(3817)


Problem 13

Replace every copy of x with 5. Then follow the order of operations PEMDAS to simplify.

f(x) = -13x^2-7x^4 - 11x
f(5) = -13(5)^2-7(5)^4 - 11(5)
f(5) = -13(25)-7(625) - 11(5)
f(5) = -325-4375-55
f(5) = -4755

Answer: -4755

====================================================

Problem 14

Use the same idea to compute g(7) like so
g(x) = -15x^2+14x-10
g(7) = -15(7)^2+14(7)-10
g(7) = -647

This means
g(g(7))
is the same as
g(-647)

And,
g(x) = -15x^2+14x-10
g(-647) = -15(-647)^2+14(-647)-10
g(-647) = -6,288,203

Therefore, g(g(7)) = -6,288,203
For more information, search out "function composition" or "composite functions".

Answer: -6,288,203

Answer by josgarithmetic(39618)

No help on #13 needed.

#14, just try forming the composition, and substitute 7. Evaluate. That's all. In fact on this one, you can first find g(7); and then plug the value found into g(x) again and evaluate.